50.0 g of water at 22 °C is mixed with 125 g of water initially at 36 °C. What is the final temperature of the water after mixing, assuming no heat is lost to the surroundings?
Answer #1

Let the final temperature be T , for hot water temperature change = 36-T ,

for cold water temperature change = T-22

Heat lost by hot water = Heat gained by cold water

specific heat of water x mass of hot water x temperature change = specific heat of water x mass of cold water x temp change

4.184 J/gC x 125 g x ( 36-T) = 4.184 J/gC x 50g x ( T-22)

125 ( 36-T) = 50 ( T-22)

4500 - 125T = 50T - 1100

T = 32 C

Thus final temp is 32 C